The average of 5 consecutive odd numbers V,W,X,Y and Z is 41.What is the product of V and Z?
1
Expert's answer
2011-04-07T05:54:37-0400
If the numbers are consecutive and odd,
W = V + 2
X =& V + 4
Y =& V + 6
Z =& V + 8
The mean is (V+ (V+2) + (V+4) + (V+6) + (V+8))/5& = V + 4 = 41.
V = 41 - 4& = 37,&
Z = 37 + 8 = 45.
The product of V and Z is 37*45 = 1665.